Circular No. 7474

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000da IN UGC 5 A. B. Aazami and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126) report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Aug. 4.5 (mag ∼ 17.8) and 5.4 UT (mag ∼ 17.5). The new object is located at α = 0^{h}03^{m}06^{s}.52, δ = −1^{o}54′41′′.8 (equinox 2000.0), which is 12′′.8 east and 9′′.3 north of the nucleus of UGC 5. A KAIT image of the same field on July 24.5 UT showed nothing at the position of SN 2000da (limiting mag ∼ 19.0).

SUPERNOVA 2000cz IN IC 1535 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics; and P. Garnavich, University of Notre Dame, report that a spectrum of SN 2000cz (cf. IAUC 7470), obtained by M. Calkins on Aug. 3.46 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova at or slightly after maximum light. The spectrum exhibits a flat continuum with P-Cyg Balmer lines, He I, and Fe II. Adopting the NASA/IPAC Extragalactic Database recession velocity of 5231 km/s for the host galaxy yields an expansion velocity for the supernova of 4900 km/s, based on the Hα profile.

D. Rich, Hampden, ME, reports a prediscovery observation of SN 2000cz at mag ∼ 17.7 in an image taken with a 0.26-m telescope on July 24.24 UT.

COMET C/1999 S4 (LINEAR) Further to IAUC 7472, M. Kidger reports that the limiting magnitude

for a 5σ point source detection in the Isaac Newton Telescope images taken on Aug. 1.9 UT is R = 22.0. He adds that the continued presence of a well-defined leading edge or point to the coma suggests that an unresolved fragment of the nucleus in this position continues to release dust. However, this structure has faded considerably since the disruption occurred. Measurements of the coma brightness distribution indicate that the maximum of light is ∼ 75′′ from the leading border of the coma.

Visual m_{1} estimates: July 31.85 UT, 7.7 (M. Lehký, Hradec Králové, Czech Republic, 25×100 binoculars); Aug. 1.85, 8.5 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); 2.35, 9.0 (D. A. J. Seargent, The Entrance, N.S.W., 25×100 binoculars).

2000 August 5 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT Carl W. Hergenrother